I have 80 students 2 buses that cost 960 a piece and lunch is 10 dollars per student. What is total cost per student in proper math formula?
Found 2 solutions by Theo, MathTherapy:
Answer by Theo(13342)

total cost is equal to 2 * 960 + 10 * 80.
cost per student would be that / 80.
your equation becomes total cost per student = (2 * 960 + 10 * 80) / 80.
that comes out to be 34 dollars per student.

multiply that by 80 to get a total cost of 2720.
2 * 960 + 10 * 80 = 1920 + 800 = 2720.
the equation is confirmed to be good.

Answer by MathTherapy(10551)

I have 80 students 2 buses that cost 960 a piece and lunch is 10 dollars per student. What is total cost per student in proper math formula?
Since an EQUATION is required, then let total cost to EACH student be C
Then we get: